OpenStudy (anonymous):

6x(5+7)/4-7x2+3x3 i am using PEMDAS and coming up with -5 but the right answer according to the TEAS Test book that I am studying is 13. Please explain why that this is the right answer.

OpenStudy (matt101):

6x(5+7)/4-7x2+3x3 = 6x(12)/4-14+9 =18-14+9 =13

OpenStudy (anonymous):

So if subtraction comes first instead of the addition you do it ?

OpenStudy (matt101):

It doesn't matter which order you do addition and subtraction since they are actually the same function (subtraction is like adding a negative number). In your equation above, you can do the addition step first, though note you are adding 9 to -14, not 14. You will get the same answer. Side note: similarly, multiplication and division are similar functions since division is like multiplying by a reciprocal.

OpenStudy (anonymous):

Thanks for your help math is not my best subject and this test is really import to me. PS. Now I feel kind of dumb because now that you explained all of that I am starting to remember. :)

OpenStudy (matt101):

No problem - I'm happy to help :)

OpenStudy (anonymous):

Girl05 P E D M A S is your order of operations

OpenStudy (anonymous):

so its different from bedmas

OpenStudy (anonymous):

( ) exponents / x + -

OpenStudy (anonymous):

also know as Please Excuse My Dear Aunt Sally

OpenStudy (anonymous):

but bedmas b=() e=exponents d=division m=multiplication a=addition s=subtraction so in pedmas what does the p stand for

OpenStudy (matt101):

BEDMAS and PEMDAS are exactly the same thing (B = brackets, P = parentheses). They rearrange the letters a bit to make fancy acronyms, but it doesn't make a difference. D and M can be performed in any order, and A and S can be performed in any order because, as I mentioned above, they are essentially the same operation.

OpenStudy (anonymous):

matt101 thanks i didnt get the difference between p and b but now youve explained it to me so thanks

OpenStudy (anonymous):

I didn't know the b and p are the same thing.

OpenStudy (anonymous):

neither did i

OpenStudy (matt101):

Haha well I'm glad I was able to teach you both something new